How do I find the missing sides of the triangles when each trianlge measues 30,60,90 degrees and the leg is 8
Thank you
: How do I find the missing sides of the triangles when each trianlge measues 30,60,90 degrees and the leg is 8
Thank you

Answer by stanbon(18788)

How do I find the missing sides of the triangles when each trianlge measues 30,60,90 degrees and the leg is 8
----------------------------------
That depends on whether the 8 is opposite the 30 or the 60 degree angle.
--------------------
If it is opposite the 30 degree angle, it is half the hypotenuse, so
the hypotenuse would have to be 16 and the 3rd side would have to be
sqrt[16^2 - 8^2]
----------------------
If is is opposite the 60 degree angle, the side opposite the 30 is "x"
and the hypotenuse is 2x.
Then (2x)^2 = x^2 + 8^2
3x^2 = 64
x^2 = 64/3
x = 8/sqrt(3)
2x = 16/sqrt(3)
=====================
Cheers,
Stan H.